OpenStudy (anonymous):

integral of x/((x^2 +2)^2) from 0 to infinity

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

how did you do it?

OpenStudy (anonymous):

First, what you need to do is notice that this is an improper integral then rewrite as \[ \lim_{a \rightarrow \infty} \int\limits_{0}^{\infty} x/(x^2+2)^2 dx\] now do u substitution \[u = x^2+2\] \[du =2x dx\] \[du/2 = x dx\] change the integral with respect to u \[\lim_{a \rightarrow \infty} 1/2\int\limits_{2}^{a} 1/u^2 du\] don't forget to alter the integration bounds 0 and a, for a, this doesn't matter as this is representative of infinity. for 0, note that u=x^2 + 2 so the new lower bound is 0^2 +2 = 2 integrate \[\lim_{a \rightarrow \infty} 1/2\ [-1/a + 1/2] = 1/4\] note that -1/a where a can be "very large" as it goes to infinity makes -1/a go to zero I'm new to this site so it took me awhile to get the fancy LaTeX format up. sorry for the delay

OpenStudy (anonymous):

oops, in the first equation, note that infinity is supposed to be subbed with a